# SCR²-ST: Combine Single Cell with Spatial Transcriptomics for Efficient Active Sampling via Reinforcement Learning

### v1

@@ 1 Introduction
Spatial transcriptomics (ST) provides a new perspective for studying the relationship between pathological tissue structures and their spatial gene expression patterns (burgess2019spatial; asp2019spatiotemporal; asp2020spatially). However, acquiring ST data remains relatively expensive (choe2023advances), which together pose challenges for large-scale data collection in practice (he2020integrating; zhu2025asign).
Histology features exhibit strong correlations with gene expression patterns (badea2020identifying), providing a foundation for image-based gene expression prediction (he2020integrating; zhu2025computer). Deep learning methods have begun leveraging histology images to infer ST expression profiles of each tissue slide (xie2024spatially; yang2023exemplar; zhu2025img2st; zhu2025magnet). Representative approaches include regression-based ST-Net (he2020integrating), HisToGene (pang2021leveraging), and EGN (yang2023exemplar), which directly predict expression values from local image appearance; and retrieval-based vision–omics contrastive learning methods, such as BLEEP (xie2024spatially) and mlxSTExp (min2024multimodal).
However, traditional fixed-grid sampling inevitably acquires many spatially adjacent regions with highly similar morphology, leading to substantial molecular redundancy and reduced biological diversity. Its non-selective nature also results in the inclusion of biologically uninformative areas (schroeder2025scaling; grases2025practical). Consequently, the effective information density of the dataset is low, causing a mismatch between sequencing cost and informative yield, thus constraining the performance and scalability of image-based ST prediction methods.
The limited availability of ST data motivates the integration of external biological knowledge to compensate for inherent constraints in coverage and data quality. In particular, the single-cell sequencing field provides substantially richer priors, supported by large-scale datasets (regev2017human) and powerful foundation models (cui2024scgpt), with sample sizes typically exceeding ST by more than an order of magnitude (svensson2018exponential). Single-cell profiles resolve cellular types, states, and regulatory programs (cao2019single; stuart2019comprehensive), offering mechanistic insight into gene expression variation across tissues. Incorporating such fine-grained priors into ST analysis introduces valuable structural guidance and biological constraints, helping mitigate challenges related to limited sampling.
@@ 2.2.1 Policy Network for Active Sampling
Prior to active sampling, we perform dense visual feature extraction on tissue sections. Specifically, we uniformly partition WSIs patches and employ the pre-trained UNI (chen2024uni) to extract visual embeddings $\{e_{i}\}_{i=1}^{N}$ for each patch, along with their corresponding spatial coordinates $\{(x_{i},y_{i},w_{i})\}_{i=1}^{N}$, where $w_{i}$ denotes the slide identifier. Based on these embeddings, we construct a lightweight policy network $\pi_{\theta}(\cdot)$ that outputs a sampling priority score for each candidate location as
@@ 2.2.2 Multi-Objective Reward Design
After obtaining sample set $S_{t}$ at round $t$, we evaluate sampling quality and construct multi-objective reward signals to update the policy network. We extract ST expression embeddings $\{\mathbf{z}_{i}\}_{i\in S_{t}}$ by pretrained scGPT (cui2023scGPT) for sampled locations and reference embeddings $\{\mathbf{q}_{j}\}_{j=1}^{M}$ from external single-cell data. The reward function measures sampling quality from two complementary perspectives: biological diversity and spatial uniformity.
@@ 2.3.2 Regression Path and Training Objective
We adopt DenseNet-121 (huang2017densely) pre-trained on ImageNet as the visual encoder to capture histomorphological patterns. Given an input patch, the encoder produces a compact feature vector through global average pooling. A two-layer MLP then decodes the visual features into gene expression predictions $\hat{y}$, forming the direct regression path.
@@ 3 Data and Experiments
Datasets and Preprocessing. We evaluate all methods using three public ST datasets, including HER2 (andersson2021spatial), Breast Cancer (he2020integrating), and Kidney (lake2023atlas). For each spot, we cropped a $224\times 224$ pixel patch centered on spatial coordinates as model input. We selected top 300 genes with highest average variance as prediction targets. Following BLEEP (xie2024spatially), we applied a $\log(1+x)$ transformation on raw readouts. For the external single-cell datasets, we use two million cells from (lake2025cellular) as the reference for the Kidney dataset, and around three million cells from (chen2025highly; reed2024single; klughammer2024multi) as the reference for the Breast Cancer and HER2 datasets. A detailed profile of datasets is provided in Appendix A.
Baseline. We compared our model against SOTA methods, including regression-based models ST-Net (he2020integrating), EGN (yang2023exemplar), HisToGene (pang2021leveraging), His2ST (zeng2022spatial), and TRIPLEX (chung2024accurate), and retrieval-based models BLEEP (xie2024spatially) and mclSTExp (min2024multimodal). All methods were trained and evaluated under consistent experimental settings to ensure fair comparison. To validate our sampling strategy, we select Monte Carlo random sampling, uncertainty-based sampling (safaei2024entropic), and diversity-driven sampling (zhdanov2019diverse) for comparison.
@@ 4.2 Empirical Validation on Gene Expression Prediction
We conducted four-fold cross-validation at the sample level to validate SCR2Net against SOTAs. Table 1 summarizes quantitative comparisons across different cohorts. Our SCR2Net outperforms existing methods in almost all metrics, achieving the lowest MSE and MAE $\downarrow$ as well as the highest PCC on most datasets. For example, on the Kidney dataset, SCR2Net improves PCC by a clear margin to 0.2391, compared with prior baselines of 0.2020. Furthermore, as illustrated in Figure 4, SCR2Net maintains strong predictive robustness under varying sampling budgets with our SCRL sampling strategy. The performance gap is pronounced under low sampling ratios (10–25%), where other approaches suffer degradation due to sparse tissue coverage. In contrast, SCR2Net mitigates this by acquiring informative tissue regions and leveraging retrieval-based auxiliary priors, resulting in a more reliable predictive performance with reduced sequencing costs.
Figure 4: Comparison of different sampling ratios (10-75%) across datasets. Our SCR2Net consistently achieves lower MSE and higher PCC, especially under low sampling budgets.
@@ 4.3 Ablation Study
Functional Blocks in SCR2Net. As shown in Table 2, removing the Retrieval Reference Module leads to an increase in MSE $\downarrow$ from 0.7038 to 0.7460 and a decrease in PCC $\uparrow$ from 0.2391 to 0.1851 on the Kidney dataset, which indicates its effectiveness in providing reference priors by incorporating similar spots. Meanwhile, the majority cell type filtering mechanism suppresses interference from noisy references by excluding low-quality retrieved spots, thereby enhancing the biological relevance and quality of retrieval references.
@@ Appendix A Additional Details of Dataset
For each spot location, we extracted a $224\times 224$ pixel histology patch centered on its spatial coordinate as model input. To construct the prediction targets, we selected the top 300 genes with the highest variance in expression within each dataset. Following BLEEP [xie2024spatially], we applied a $\log(1+x)$ transformation to the raw count matrices to alleviate the heavy-tailed distribution characteristic of ST expression data [he2020integrating]. The dataset-specific selected genes are visualized in Appendix Figure 5.
@@ Appendix B Additional Implementation Details
To evaluate the effectiveness of our sampling strategy, we compare it against three representative baselines: Monte Carlo random sampling, uncertainty-based sampling [safaei2024entropic], and diversity-driven sampling [zhdanov2019diverse].
@@
Diversity-driven sampling. This method encourages sample diversity based on feature similarity. We first extract 1024-dimensional visual features using the vision encoder during training. These features are standardized and reduced to 128 dimensions via PCA, followed by clustering using DBSCAN. To avoid insufficient cluster granularity, we incorporate a dynamic adjustment mechanism: the minimum cluster count is adaptively set to $\sqrt{N}/5$, where NN denotes the total number of samples. If DBSCAN yields too few clusters, we automatically switch to KMeans to enforce the desired number of clusters. During sampling, patches are drawn uniformly from each cluster to ensure diverse coverage of tissue regions.
@@ Appendix C Additional Experimental Results
Due to space limitations in the main manuscript, we provide additional experimental results in this appendix, including the comparison of different sampling ratios across datasets (Figure 4), the comparison of sampling strategies evaluated on mlSTExp and TRIPLEX benchmarks (Figure 7), and an ablation study on the reward function design (Figure 6).

### v2

@@ 1 Introduction
Spatial transcriptomics (ST) provides a new perspective for studying the relationship between pathological tissue structures and their spatial gene expression patterns (Burgess, 2019; Asp et al., 2019, 2020). However, acquiring ST data remains relatively expensive (Choe et al., 2023), which together pose challenges for large-scale data collection in practice (He et al., 2020; Zhu et al., 2025d).
Histology features exhibit strong correlations with gene expression patterns (Badea and Stănescu, 2020), providing a foundation for image-based gene expression prediction (He et al., 2020; Zhu et al., 2025a). Deep learning methods have begun leveraging histology images to infer ST expression profiles of each tissue slide (Xie et al., 2024; Yang et al., 2023; Zhu et al., 2025b, c). Representative approaches include regression-based ST-Net (He et al., 2020), HisToGene (Pang et al., 2021), and EGN (Yang et al., 2023), which directly predict expression values from local image appearance; and retrieval-based vision–omics contrastive learning methods, such as BLEEP (Xie et al., 2024) and mlxSTExp (Min et al., 2024).
However, traditional fixed-grid sampling inevitably acquires many spatially adjacent regions with highly similar morphology, leading to substantial molecular redundancy and reduced biological diversity. Its non-selective nature also results in the inclusion of biologically uninformative areas (Schroeder et al., 2025; Grases and Porta-Pardo, 2025). Consequently, the effective information density of the dataset is low, causing a mismatch between sequencing cost and informative yield, thus constraining the performance and scalability of image-based ST prediction methods.
The limited availability of ST data motivates the integration of external biological knowledge to compensate for inherent constraints in coverage and data quality. In particular, the single-cell sequencing field provides substantially richer priors, supported by large-scale datasets (Regev et al., 2017) and powerful foundation models (Cui et al., 2024), with sample sizes typically exceeding ST by more than an order of magnitude (Svensson et al., 2018). Single-cell profiles resolve cellular types, states, and regulatory programs (Cao et al., 2019; Stuart et al., 2019), offering mechanistic insight into gene expression variation across tissues. Incorporating such fine-grained priors into ST analysis introduces valuable structural guidance and biological constraints, helping mitigate challenges related to limited sampling.
@@ 2.2.1 Policy Network for Active Sampling
Prior to active sampling, we perform dense visual feature extraction on tissue sections. Specifically, we uniformly partition WSIs patches and employ the pre-trained UNI (Chen et al., 2024) to extract visual embeddings $\{e_{i}\}_{i=1}^{N}$ for each patch, along with their corresponding spatial coordinates $\{(x_{i},y_{i},w_{i})\}_{i=1}^{N}$, where $w_{i}$ denotes the slide identifier. Based on these embeddings, we construct a lightweight policy network $\pi_{\theta}(\cdot)$ that outputs a sampling priority score for each candidate location as
@@ 2.2.2 Multi-Objective Reward Design
After obtaining sample set $S_{t}$ at round $t$, we evaluate sampling quality and construct multi-objective reward signals to update the policy network. We extract ST expression embeddings $\{\mathbf{z}_{i}\}_{i\in S_{t}}$ by pretrained scGPT (Cui et al., 2023) for sampled locations and reference embeddings $\{\mathbf{q}_{j}\}_{j=1}^{M}$ from external single-cell data. The reward function measures sampling quality from two complementary perspectives: biological diversity and spatial uniformity.
@@ 2.3.2 Regression Path and Training Objective
We adopt DenseNet-121 (Huang et al., 2017) pre-trained on ImageNet as the visual encoder to capture histomorphological patterns. Given an input patch, the encoder produces a compact feature vector through global average pooling. A two-layer MLP then decodes the visual features into gene expression predictions $\hat{y}$, forming the direct regression path.
@@ 3 Data and Experiments
Datasets and Preprocessing. We evaluate all methods using three public ST datasets, including HER2 (Andersson et al., 2021), Breast Cancer (He et al., 2020), and Kidney (Lake et al., 2023). For each spot, we cropped a $224\times 224$ pixel patch centered on spatial coordinates as model input. We selected top 300 genes with highest average variance as prediction targets. Following BLEEP (Xie et al., 2024), we applied a $\log(1+x)$ transformation on raw readouts. For the external single-cell datasets, we use two million cells from Lake et al. (2025) as the reference for the Kidney dataset, and around three million cells from Chen et al. (2025); Reed et al. (2024); Klughammer et al. (2024) as the reference for the Breast Cancer and HER2 datasets. A detailed profile of datasets is provided in Appendix A.
Baseline. We compared our model against SOTA methods, including regression-based models ST-Net He et al. (2020), EGN Yang et al. (2023), HisToGene Pang et al. (2021), His2ST Zeng et al. (2022), and TRIPLEX Chung et al. (2024), and retrieval-based models BLEEP Xie et al. (2024) and mclSTExp Min et al. (2024). All methods were trained and evaluated under consistent experimental settings to ensure fair comparison. To validate our sampling strategy, we select Monte Carlo random sampling, uncertainty-based sampling Safaei et al. (2024), and diversity-driven sampling Zhdanov (2019) for comparison.
@@ 4.1 Active Sampling under Budget Constraints
Figure 4: Comparison of different sampling ratios (10-75%). Our SCR2Net consistently achieves better metrics, especially under low sampling budget scenarios.
@@ 4.2 Empirical Validation on Gene Expression Prediction
We conducted four-fold cross-validation at the sample level to validate SCR2Net against SOTAs. Table 1 summarizes quantitative comparisons across different cohorts. Our SCR2Net outperforms existing methods in almost all metrics, achieving the lowest MSE and MAE $\downarrow$ as well as the highest PCC on most datasets. For example, on the Kidney dataset, SCR2Net improves PCC by a clear margin to 0.2391, compared with prior baselines of 0.2020. Furthermore, as illustrated in Figure 4 in Appendix C, SCR2Net maintains strong predictive robustness under varying sampling budgets with our SCRL sampling strategy. The performance gap is pronounced under low sampling ratios (10–25%), where other approaches suffer degradation due to sparse tissue coverage. In contrast, SCR2Net mitigates this by acquiring informative tissue regions and leveraging retrieval-based auxiliary priors, resulting in a more reliable predictive performance with reduced sequencing costs.
@@ 4.3 Ablation Study
Table 2: Stability of random initializations. We reported the performance comparison of our SCRL and random sampling under different random seeds, indicating that our sampling strategy achieves stable performance across different initializations.
Model Ratio SCRL (Ours) Random
[…]
Functional Blocks in SCR2Net. As shown in Table 3, removing Retrieval Reference Module leads to an increase in MSE from 0.7038 to 0.7460 and a decrease in PCC $\uparrow$ from 0.2391 to 0.1851 on the Kidney dataset, which indicates its effectiveness in providing reference priors by incorporating similar spots. Meanwhile, majority cell type filtering mechanism suppresses interference from noisy references by excluding low-quality retrieved spots.
Figure 5: Sensitivity analysis of reward weight configurations under different sampling ratios. Each curve represents a distinct reward weight setting, with random sampling shown as a baseline. When reward terms are placed at comparable scales, both models exhibit consistent performance trends as the sampling ratio increases. In contrast, configurations dominated by a single reward term result in clear performance degradation, highlighting the necessity of balancing biological diversity and spatial coverage.
Sensitivity analysis of reward weight. We conduct a sensitivity analysis of the reward weight configurations by systematically controlling the weights of different reward terms at different scales to evaluate how scale differences affect sampling behavior and downstream prediction performance. Specifically, we construct multiple weight configurations, where some place all reward terms within a comparable scale (e.g., the original setting and nearby configurations), while others assign certain reward terms to significantly different orders of magnitude. As shown in Figure 5 and Figure 8, only when the reward terms are adjusted to comparable scales does the model maintain good and consistent performance; when a single reward term dominates or random sampling is adopted, the performance degrades noticeably. These results suggest that the effectiveness of the proposed reward design does not rely on precise weight tuning, but rather on maintaining a balanced contribution between biological diversity and spatial coverage.
Stability under random initializations. We evaluate the stability of the proposed adaptive sampling strategy by repeating experiments with different random seeds under multiple sampling ratios. As shown in Table 2, SCRL consistently achieves stable performance with small variance across runs and outperforms random sampling, indicating that the learned sampling behavior is robust to random initialization.
Limitations and Challenges. Despite these gains, overall accuracy remains bounded by intrinsic challenges of spatial transcriptomics rather than the sampling strategy alone. The most challenging regions for both sampling and prediction are typically highly heterogeneous or transitional tissue areas, as well as regions containing rare cell populations, which are often clinically important. In addition, morphologically similar regions may exhibit distinct molecular profiles, fundamentally limiting the predictive power of image-based models. Technical noise and spot-level signal averaging inherent to ST measurements further introduce uncertainty.
@@ Appendix A Additional Details of Dataset
For each spot location, we extracted a $224\times 224$ pixel histology patch centered on its spatial coordinate as model input. To construct the prediction targets, we selected the top 300 genes with the highest variance in expression within each dataset. Following BLEEP Xie et al. [2024], we applied a $\log(1+x)$ transformation to the raw count matrices to alleviate the heavy-tailed distribution characteristic of ST expression data He et al. [2020]. The dataset-specific selected genes are visualized in Appendix Figure 6.
@@ Appendix B Additional Implementation Details
To evaluate the effectiveness of our sampling strategy, we compare it against three representative baselines: Monte Carlo random sampling, uncertainty-based sampling Safaei et al. [2024], and diversity-driven sampling Zhdanov [2019].
@@
Table 4: Summary of computational cost. We summarize the parameter size, running time per epoch, GPU memory consumption, and number of training samples of our model on Breast Cancer dataset.
Module Parameter Time/epoch GPU Mem Samples
SCR2Net 33.82 M 85.70 s 35.13 GB 27,171
SCRL 0.6 M 50.54 s 3,099,206
@@
Diversity-driven sampling. This method encourages sample diversity based on feature similarity. We first extract 1024-dimensional visual features using the vision encoder during training. These features are standardized and reduced to 128 dimensions via PCA, followed by clustering using DBSCAN. To avoid insufficient cluster granularity, we incorporate a dynamic adjustment mechanism: the minimum cluster count is adaptively set to $\sqrt{N}/5$, where $N$ denotes the total number of samples. If DBSCAN yields too few clusters, we automatically switch to KMeans to enforce the desired number of clusters. During sampling, patches are drawn uniformly from each cluster to ensure diverse coverage of tissue regions.
Figure 9: Spatial distribution of sampled spots under 10% sampling. Red circles indicate selected spots on whole-slide images. Entropy-based sampling produces concentrated regions, while diversity-based sampling wastes samples in low-density background areas. Our method (SCRL) achieves more informative and balanced sampling by incorporating spatial and biological cues.
@@
Figure 10: Visualization of predicted spatial expression distribution of cancer-related gene RPS3 on WSIs.
@@ Appendix C Additional Experimental Results
Due to space limitations in the main manuscript, we provide additional experimental results in this appendix, including the qualitative analysis of low-budget sampling (Figure 9), the comparison of sampling strategies evaluated on mlSTExp and TRIPLEX benchmarks (Figure 7), and an ablation study on the reward function design (Figure 8).
Figure 11: Visualization of predicted spatial expression distribution of cancer-related gene RPL19 on WSIs.
Qualitative analysis of low-budget sampling. Figure 8 further provides a qualitative comparison of different sampling strategies by visualizing the spatial distribution of sampled spots on whole-slide images under a low sampling budget (10%). As shown in the figure, entropy-based sampling tends to select spots from relatively concentrated regions, resulting in limited spatial coverage of the tissue. In contrast, diversity-based sampling distributes samples more broadly, but often allocates a substantial portion of the limited sampling budget to low tissue-density or background regions, leading to inefficient sampling. By jointly incorporating spatial relationships and biological information, our method achieves a more balanced sampling pattern that focuses on informative tissue regions while maintaining adequate spatial coverage. This behavior explains the superior performance of our approach under low-budget sampling settings, where effective utilization of limited sampling resources is critical.
Cancer-related Gene Expression Prediction. We further selected two genes that are highly relevant to breast cancer, RPL19 and RPS3. RPL19 is frequently overexpressed in breast cancer tissues, making it a commonly used reference and prognostic-related gene in breast cancer studies [hong2014ribosomal]. RPS3’s dysregulation is associated with enhanced tumor aggressiveness and poor prognosis in breast cancer [ono2017ribosomal]. We visualize the predicted spatial expression patterns of these two genes on whole-slide images in Figure 11 and Figure 10, demonstrating that the proposed model captures spatially coherent and biologically meaningful expression distributions, thereby highlighting its interpretability at clinically relevant molecular levels.
Appendix D Practical Considerations and Future Directions
From a practical perspective, the reproducibility and applicability of the proposed framework in new experimental settings can be considered from both the active sampling strategy (SCRL) and the gene expression prediction model (SCR2Net). The SCRL sampler is designed to operate with any amount of existing ST data and does not require full-coverage tissue sampling at initialization. As summarized in Table 2 and discussed in Section 4.3, SCRL exhibits consistent and stable performance across different initializations, indicating robustness to the choice of starting point. This behavior is attributed to the formulation of SCRL as a sequential decision-making process, where the sampler is optimized online based on the information observed so far and iteratively selects tissue locations to query until the predefined sampling budget is reached.
For SCR2Net, we leverage pretrained weights for the vision encoder to enhance representation learning and improve generalization across datasets. However, in real-world applications, batch effects and inter-laboratory variations in tissue preparation and sequencing protocols often introduce a domain gap between the training data and unseen experimental settings. Consequently, fine-tuning the model with a small amount of data from the target laboratory is typically required, which is a common and practical assumption in spatial transcriptomics modeling.
As more data become available in future studies, both the prediction model and the active sampling strategy can be further optimized, which is expected to improve the robustness, scalability, and transferability of the overall framework across diverse experimental conditions.